$SPAD/src/input radff.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool radff.output
)set message test on
)set message auto off
)clear all
--S 1 of 27
P0 := UP(x, INT)
--R
--R
--R   (1)  UnivariatePolynomial(x,Integer)
--R                                                                 Type: Domain
--E 1
--S 2 of 27
P1 := UP(y, FRAC P0)
--R
--R
--R   (2)  UnivariatePolynomial(y,Fraction(UnivariatePolynomial(x,Integer)))
--R                                                                 Type: Domain
--E 2

———-

curve given by x^{2}0 + y^{2}0 = 1 — * —

--S 3 of 27
R := RADFF(INT, P0, P1, 1 - x**20, 20)
--R
--R
--R   (3)
--R  RadicalFunctionField(Integer,UnivariatePolynomial(x,Integer),UnivariatePolyno
--R  mial(y,Fraction(UnivariatePolynomial(x,Integer))),(-(x^20))+1,20)
--R                                                                 Type: Domain
--E 3
--S 4 of 27
definingPolynomial()$R
--R
--R

--R         20    20
--R   (4)  y   + x   - 1
--R      Type: UnivariatePolynomial(y,Fraction(UnivariatePolynomial(x,Integer)))
--E 4
--S 5 of 27
genus()$R
--R
--R
--R   (5)  171
--R                                                     Type: NonNegativeInteger
--E 5
--S 6 of 27
rank()$R
--R
--R
--R   (6)  20
--R                                                        Type: PositiveInteger
--E 6
--S 7 of 27
numberOfComponents()$R
--R
--R
--R   (7)  1
--R                                                     Type: NonNegativeInteger
--E 7
--S 8 of 27
integralBasisAtInfinity()$R
--R
--R
--R   (8)
--R       1     1  2   1  3   1  4   1  5   1  6   1  7   1  8   1  9   1   10
--R   [1, - y, -- y , -- y , -- y , -- y , -- y , -- y , -- y , -- y , --- y  ,
--R       x     2      3      4      5      6      7      8      9      10
--R            x      x      x      x      x      x      x      x      x
--R     1   11   1   12   1   13   1   14   1   15   1   16   1   17   1   18
--R    --- y  , --- y  , --- y  , --- y  , --- y  , --- y  , --- y  , --- y  ,
--R     11       12       13       14       15       16       17       18
--R    x        x        x        x        x        x        x        x
--R     1   19
--R    --- y  ]
--R     19
--R    x
--RType: Vector(RadicalFunctionField(Integer,UnivariatePolynomial(x,Integer),UnivariatePolynomial(y,Fraction(UnivariatePolynomial(x,Integer))),(-(x^20))+1,20))
--E 8
--S 9 of 27

branchPoint?(0)$R
--R
--R
--R   (9)  false
--R                                                                Type: Boolean
--E 9
--S 10 of 27
branchPoint?(1)$R
--R
--R
--R   (10)  true
--R                                                                Type: Boolean
--E 10
--S 11 of 27
y := generator()$R
--R
--R
--R   (11)  y
--RType: RadicalFunctionField(Integer,UnivariatePolynomial(x,Integer),UnivariatePolynomial(y,Fraction(UnivariatePolynomial(x,Integer))),(-(x^20))+1,20)
--E 11
--S 12 of 27
norm y
--R
--R
--R          20
--R   (12)  x   - 1
--R                              Type: Fraction(UnivariatePolynomial(x,Integer))
--E 12
--S 13 of 27
trace y
--R
--R
--R   (13)  0
--R                              Type: Fraction(UnivariatePolynomial(x,Integer))
--E 13

———-

curve given by y^{4} = 2x^{2} — * —

--S 14 of 27
R2 := RADFF(INT, P0, P1, 2 * x**2, 4)
--R
--R
--R   (14)

--R  RadicalFunctionField(Integer,UnivariatePolynomial(x,Integer),UnivariatePolyno
--R  mial(y,Fraction(UnivariatePolynomial(x,Integer))),2*x^2,4)
--R                                                                 Type: Domain
--E 14
--S 15 of 27
definingPolynomial()$R2
--R
--R
--R          4     2
--R   (15)  y  - 2x
--R      Type: UnivariatePolynomial(y,Fraction(UnivariatePolynomial(x,Integer)))
--E 15
--S 16 of 27
rank()$R2
--R
--R
--R   (16)  4
--R                                                        Type: PositiveInteger
--E 16
--S 17 of 27
absolutelyIrreducible?()$R2
--R
--R
--R   (17)  false
--R                                                                Type: Boolean
--E 17
--S 18 of 27
numberOfComponents()$R2
--R
--R
--R   (18)  2
--R                                                     Type: NonNegativeInteger
--E 18
--S 19 of 27
genus()$R2
--R
--R
--R   (19)  0
--R                                                     Type: NonNegativeInteger
--E 19
--S 20 of 27
integralBasis()$R2
--R
--R

--R              1  2 1  3
--R   (20)  [1,y,- y ,- y ]
--R              x    x
--RType: Vector(RadicalFunctionField(Integer,UnivariatePolynomial(x,Integer),UnivariatePolynomial(y,Fraction(UnivariatePolynomial(x,Integer))),2*x^2,4))
--E 20
--S 21 of 27
y := generator()$R2
--R
--R
--R   (21)  y
--RType: RadicalFunctionField(Integer,UnivariatePolynomial(x,Integer),UnivariatePolynomial(y,Fraction(UnivariatePolynomial(x,Integer))),2*x^2,4)
--E 21
--S 22 of 27
integralCoordinates(y**3)
--R
--R
--R   (22)  [num= [0,0,0,x],den= 1]
--RType: Record(num: Vector(UnivariatePolynomial(x,Integer)),den: UnivariatePolynomial(x,Integer))
--E 23
--S 24 of 27
integralRepresents(%.num, %.den)$R2
--R
--R
--R          3
--R   (23)  y
--RType: RadicalFunctionField(Integer,UnivariatePolynomial(x,Integer),UnivariatePolynomial(y,Fraction(UnivariatePolynomial(x,Integer))),2*x^2,4)
--E 24
--S 25 of 27
norm y
--R
--R
--R             2
--R   (24)  - 2x
--R                              Type: Fraction(UnivariatePolynomial(x,Integer))
--E 25
--S 26 of 27
trace y
--R
--R
--R   (25)  0
--R                              Type: Fraction(UnivariatePolynomial(x,Integer))
--E 26
--S 27 of 27
regularRepresentation y

--R
--R
--R         + 0   1  0  0+
--R         |            |
--R         | 0   0  1  0|
--R   (26)  |            |
--R         | 0   0  0  1|
--R         |            |
--R         |  2         |
--R         +2x   0  0  0+
--R                      Type: Matrix(Fraction(UnivariatePolynomial(x,Integer)))
--E 27
)spool
)lisp (bye)

———-

References^{[1] nothing}

9